OpenStudy (anonymous):

Let c be any constant number. Which of the following will always be perpendicular to -4x + y = -2? Answer y equals negative one over four times x plus c y equals four x plus c y equals one over four times x plus c y equals negative four x plus c

OpenStudy (helder_edwin):

if you have two lines with slopes m1 and m2, in order for them to be perpendicular is should be \[ \large m_1\cdot m_2=? \]

OpenStudy (anonymous):

hmm

OpenStudy (helder_edwin):

if you have the lines \[ \large y=m_1x+b_1\qquad\text{and}\qquad y=m_2x+b_2 \] with slopes m1 and m2. for them to be perpendicular m1 and m2 must satisfy \[ \large m_1\cdot m_2= \]